### v2

# CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation

Abstract
With the rapid advancement of Large Language Models (LLMs), the demand for robust instruction-following capabilities in code generation tasks has grown significantly. Code generation not only facilitates faster prototyping and automated testing, but also augments developer efficiency through improved maintainability and reusability of code. In this paper, we introduce CodeIF, the first benchmark specifically designed to assess the abilities of LLMs to adhere to task-oriented instructions within diverse code generation scenarios. CodeIF encompasses a broad range of tasks, including function synthesis, error debugging, algorithmic refactoring, and code explanation, thereby providing a comprehensive suite to evaluate model performance across varying complexity levels and programming domains. We conduct extensive experiments with LLMs, analyzing their strengths and limitations in meeting the demands of these tasks. The experimental results offer valuable insights into how well current models align with human instructions, as well as the extent to which they can generate consistent, maintainable, and contextually relevant code. Our findings not only underscore the critical role that instruction-following LLMs can play in modern software development, but also illuminate pathways for future research aimed at enhancing their adaptability, reliability, and overall effectiveness in automated code generation.  [^1] CodeIF data and code are publicly available: https://github.com/lin-rany/codeIF.
Machine Learning, ICML
1Beihang University, 2Tsinghua University, 3National University of Singapore
footnotetext: *Equal contributionfootnotetext: Corresponding Author
1 Introduction
With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation. While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge (Liu et al., 2024; Zhang et al., 2023). To drive progress in this field, a comprehensive and systematic evaluation framework is essential (Jiang et al., 2024; Zhou et al., 2023).
This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation. Built upon insights from existing evaluation sets like McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++. It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
To ensure rigorous assessment, we propose four novel evaluation metrics: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics measure models’ ability to handle multi-constraint problems from different perspectives, including full compliance, average constraint satisfaction, logical coherence, and consistency in instruction execution. By offering a structured evaluation framework, CodeIF provides valuable insights into the current state and future direction of LLM-driven code generation.
Overall, our contributions are mainly four-fold:
  1. 1.
    Innovative Benchmark: We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation. CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
  2. 2.
    Automated High-Quality Instruction Generation: Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists. This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
  3. 3.
    Novel Evaluation Metrics: We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks. These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
  4. 4.
    Extensive Evaluation and Analysis: We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels. Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
2 CodeIF
Overview:
Figure 1: The construction process of CodeIF. The first step involves the construction of constraint instructions, followed by the assembly of the dataset, and finally the construction of dependencies between instructions.
As shown in Figure 1, the construction of CodeIF includes the following steps: we first collect and refine constraint instructions from real code generation tasks, then we combine code generation tasks with LLMs and human review to create a high-quality evaluation dataset.
2.1 Building
The construction of the CodeIF dataset consists of two main phases: collection of constraint instructions and data processing. In Section  2.2, we elaborate on how we collected the constraint instructions for code generation. Following this, in Section  2.3, we describe how we combined these constraint instructions to produce the final CodeIF evaluation dataset.
2.2 Constraint Instructions Collection
The first phase of our work focuses on code generation by collecting and formulating constraint instructions to construct the CodeIF evaluation dataset. This process consists of two stages: (1) collecting and verifying constraint instructions and (2) applying them to dataset generation.
We begin with an in-depth analysis of code generation tasks based on existing benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), establishing an instruction system covering eight categories. Each category targets a specific aspect of code generation, enabling a fine-grained assessment of LLMs’ instruction-following abilities.
Within each category, we further refine constraints into atomic instructions with clear, explicit directives. These atomic constraints allow binary evaluation (“yes” or “no”), ensuring an objective and precise assessment of model compliance while minimizing subjectivity.
Our classification of atomic constraint instructions is guided by an analysis of existing code generation benchmarks and widely used code repositories. We observe that requirement distribution varies significantly across different levels—line-level, function-level, and full-solution generation. To ensure comprehensive coverage, we design atomic instruction categories that assess both high-level architectural constraints and fine-grained variable control.
To systematically evaluate LLMs’ instruction-following capabilities in code generation, we define eight constraint categories. The Global category ensures that generated code aligns with overarching specifications and holistic constraints. Structural Control focuses on adherence to control structures such as loops and conditionals, as well as data structures like sets and maps. Variable constraints assess details at the variable level, including naming conventions and initialization methods.
At different abstraction levels, Interface, Function, and Class constraints evaluate the model’s ability to construct well-defined program components. The File category tests handling of cross-file dependencies and external library calls. Finally, the Combination category presents a comprehensive challenge by integrating constraints from multiple dimensions, assessing how well models follow complex, multi-faceted instructions in realistic scenarios.
Figure 2: CodeIF Constraints Instruction Distribution
As shown in Figure 2, CodeIF’s distribution across different programming languages and categories is presented.
Through a detailed analysis of existing code generation benchmarks, we designed an evaluation system comprising 8 main categories and 50 fine-grained constraint instructions. This system not only enables a systematic assessment of LLMs’ performance in code generation but also establishes clear benchmarks for future research.
These instructions cover constraints ranging from high-level structural guidelines to fine-grained variable operations, ensuring a comprehensive evaluation. For instance, the Global category assesses a model’s ability to maintain overall structural integrity and logical coherence, while the Variable category evaluates its precision in variable naming and initialization.
By refining constraints into well-defined categories, we can precisely identify the strengths and weaknesses of different LLMs, guiding optimization efforts and advancing the field toward more effective automated code generation. Additionally, this structured evaluation approach provides a valuable reference for future research, fostering continuous innovation. The complete list of atomic constraint instructions is provided in Appendix Table 4.
2.3 Data Construction
Multi-Language and Difficulty-Differentiated Benchmark Design
To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024). These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
To further refine the evaluation, we categorize tasks into two difficulty levels: Easy and Hard. The Hard set includes longer, more intricate instruction lists, designed to rigorously test LLMs’ ability to handle complex constraints.
Automated Generation of Constraint Instructions
To construct high-quality constraint instructions, we leveraged large language models (LLMs) such as GPT-4 to automatically generate task-specific instruction lists. Initially, we curated approximately 20 detailed examples, each featuring a specific code generation task. Human experts then formulated atomic constraint instructions to ensure precise adherence during code generation.
These examples served as prompts, allowing LLMs to generate corresponding instruction lists while refining task descriptions by identifying and eliminating redundant constraints. This approach streamlined instruction-following requirements, ensuring clarity and promoting high-quality outputs.
Constructing Instruction Dependencies
Since some atomic constraints exhibit interdependencies, we further utilized LLMs to construct explicit dependency relationships within instruction sets. These dependencies were integrated into our evaluation framework, enhancing assessment depth and improving the accuracy and practicality of constraint verification.
Specifically, by analyzing the logical connections and sequential dependencies among atomic instructions, we can precisely define the execution steps required for each task. For instance, in function generation tasks, the process typically involves defining the function name first, specifying parameter types, and finally implementing the function body. These dependencies ensure coherence and consistency throughout the code generation process.
Integrating these dependencies into the evaluation system enables a more accurate assessment of a model’s ability to follow complex instructions while providing a foundation for future optimizations. Additionally, this approach helps identify potential challenges models face when handling specific instruction sequences, offering insights for further improvement. As shown in Figure 3, this example from CodeIF illustrates a specific generation task, its associated instruction list, and the corresponding instruction dependencies.
Figure 3: Specific cases of the CodeIF dataset, ’Task’ denotes the specific generation task, ’Type’ refers to the type of constraint, and ’Dependence’ indicates the prerequisite constraints for this constraint.
2.4 Data Analysis
Table 1: CodeIF dataset statistics, showing the statistical information of different difficulty classifications. Avg.Instr represents the average length of the atomic constraint instruction list.
Set Num Avg.Instr Go Python Java C++
Easy 600 11.99 127 165 176 132
Hard 600 13.80 103 183 177 137
Full 1200 12.90 230 348 353 269
CodeIF Static Analysis
Figure 4: The distribution of atomic instruction list lengths in datasets of different difficulties.
As shown in Table 1, the dataset is categorized into three difficulty levels: Easy, Hard, and an aggregated set, Full. The Easy and Hard sets each contain 600 code generation tasks, while the Full dataset combines both, totaling 1,200 tasks. These tasks span four programming languages: Go, Python, Java, and C++.
Among them, Java has the highest number of tasks (353), followed by Python (348). C++ and Go contain 269 and 230 tasks, respectively. The dataset also records the average number of instructions per task for each difficulty level. The Easy set averages 11.99 instructions per task, while the Hard set increases to 13.8, reflecting greater complexity and stricter constraints. The Full dataset, combining both, has an average instruction length of 12.9. Figure 4 illustrates the distribution of task lengths across the dataset.
Table 2: CodeIF dataset statistics information, showing the distribution of atomic restriction instruction categories under different difficulty classifications.
Set Global Structural Control Variable Interface Function Class File Combination
Easy 1638 1008 1336 427 569 544 723 953
Hard 1890 1193 1479 505 659 623 802 1142
Full 3528 2201 2815 932 1228 1167 1525 2095
Figure 5: The distribution of constraint instruction list lengths in datasets of different difficulties.
Analysis of Constraint Instruction Category Distribution
Table 2 presents the distribution of atomic instructions across different difficulty levels in the evaluation set. As shown in Figure 5, we compare the average number of instructions per category between the “Easy” and “Hard” sets. These categories include Global, Structural Control, Variable, Interface, Function, Class, File, and Combination.
Across all categories, the “Hard” dataset contains a higher average number of instructions than the “Easy” set. Notably, the Global category has the highest instruction count in both difficulty levels, averaging 2.5 instructions in the “Easy” set and exceeding 3 in the “Hard” set. This trend indicates that as task complexity increases, models face greater challenges in handling intricate constraints and detailed requirements. By analyzing instruction distribution across categories, we gain deeper insights into model performance under varying levels of difficulty, providing a foundation for further optimization.
Figure 6: Distribution of atomic instruction list lengths across difficulty levels.
Figure 6 illustrates the proportion of each atomic instruction category. The Global category holds the largest share at 22.77%, followed by the Variable category at 18.17%. This distribution highlights CodeIF’s focus on evaluating LLMs’ capabilities at both the overall structural level (e.g., Global constraints) and fine-grained variable handling (e.g., Variable constraints). This design ensures a balanced assessment of code generation, emphasizing both high-level logical coherence and precision in variable management.
3 Metrics
Ensuring that large language models (LLMs) accurately follow instructions is crucial for code generation. To precisely evaluate this capability, we introduce four novel metrics designed to assess how LLMs handle code generation tasks with multiple constraints: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics provide a comprehensive evaluation from different perspectives.
For a dataset with $m$ problems, each problem contains a set of $n_{i}$ constraints. We define CSR and SSR as follows:
Completely Satisfaction Rate (CSR)
$\text{CSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\prod_{j=1}^{n_{i}}r_{i,j}\right)$ (1)
where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied. CSR measures the proportion of problems where all constraints are fully met.
Soft Satisfaction Rate (SSR)
$\text{SSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}r_{i,j}}{n_{i}}\right)$ (2)
SSR evaluates the average proportion of constraints satisfied per problem, providing a more flexible assessment.
Rigorous Satisfaction Rate (RSR)
In code generation, some constraints depend on prior instructions, particularly in Combination constraints. To account for dependencies, we define RSR as:
$\text{RSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}\left[r_{i,j}\cdot\prod_{k\in D_{i,j}}r_{i,k}\right]}{n_{i}}\right)$ (3)
where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on. RSR ensures that models satisfy prerequisite constraints before fulfilling dependent ones.
Consistent Continuity Satisfaction Rate (CCSR)
In many code generation tasks, maintaining continuous adherence to instructions is essential. To measure this ability, we define CCSR as:
$\small\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$. CCSR evaluates a model’s consistency in following sequential instructions without errors.
4 Experiment
4.1 Baselines
We evaluate over 30 language models, including both open-source models and commercial API services.
  • Meta Llama 3 Series (Touvron et al., 2023): Includes Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct.
  • Qwen2.5 Series (Yang et al., 2024): Includes Qwen2.5-1.5B-Instruct, Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct. Additionally, it includes dedicated code generation models: Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024), Qwen2.5-Coder-3B-Instruct, Qwen2.5-Coder-7B-Instruct, Qwen2.5-Coder-14B-Instruct, and Qwen2.5-Coder-32B-Instruct.
  • Mistral Series (Jiang et al., 2023): Includes Mistral-3B, Mistral-8B, and the specialized code model Codestral-2501.
  • Microsoft Phi Series (Abdin et al., 2024): Includes Phi-3.5-Mini-128K-Instruct (3.8B) and the latest Phi-4.
  • Google Gemma Series (Team, 2024b): Includes Gemma-2-9B-It and Gemma-2-27B-It.
  • DeepSeek Series: Includes DeepSeek-Coder (Guo et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024).
  • OpenAI Series (Achiam et al., 2023): Includes GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and the latest GPT-4O-2024-11-20.
  • Google Gemini Series (Team, 2024a): Includes Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro.
  • Anthropic Series: Includes Claude-3.5-Sonnet-20241022.
4.2 Experimental Setup
The temperature coefficient is set to 0 to ensure output determinism, with a maximum generation length of 4096 tokens. All other settings follow the official default parameters for each model. Commercial API models are accessed through the latest available interface as of December 2024. Specifically, this study evaluates the Qwen2.5-Coder series, Meta’s latest Llama 3.3 architecture, and the DeepSeek-V3 large-scale dialogue model as upper-bound benchmarks. All experiments are conducted using the official API and 8 H800(80G).
4.3 Automatic Evaluation
To ensure robust evaluation, we employed both large language models (LLMs) and human experts to assess whether each model correctly followed the atomic constraint instructions. During data preprocessing, all constraints were decomposed into atomic elements, allowing the adjudication system to perform objective binary evaluations (Yes/No) instead of relying on subjective judgment.
Follow by FairEval (Wang et al., 2023a), GPT-4-1106-Preview model served as the primary automated evaluation tool (see Appendix 8.1 for the detailed prompt). Additionally, three domain experts manually annotated a stratified sample of 150 instances. Statistical analysis revealed strong agreement between LLM-based and human evaluations, with Pearson correlation coefficients of 0.87 (LLM-human) and 0.85 (inter-human), demonstrating high consistency between automated and manual assessments.
These results suggest that the atomic constraint decomposition methodology effectively reduces subjectivity, making LLM-based evaluation a reliable and scalable approach for constraint verification tasks.
4.4 Main Experiments
Table 3: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
Models CSR SSR RSR CCSR
Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
Llama-3.2-1B-Instruct 0.034 0.046 0.022 0.218 0.277 0.159 0.182 0.231 0.132 0.152 0.197 0.107
Llama-3.1-8B-Instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.340 0.370 0.444 0.295
Qwen2.5-Coder-7B-Instruct 0.142 0.198 0.087 0.514 0.590 0.438 0.453 0.533 0.373 0.390 0.463 0.318
Qwen2.5-7B-Instruct 0.153 0.201 0.104 0.535 0.599 0.471 0.475 0.546 0.405 0.416 0.479 0.353
Ministral-8B 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.490 0.371
Gemma-2-9B-It 0.171 0.210 0.131 0.573 0.642 0.504 0.513 0.587 0.440 0.445 0.508 0.383
Qwen2.5-Coder-32B-Instruct 0.365 0.422 0.307 0.736 0.767 0.704 0.679 0.723 0.635 0.634 0.669 0.599
Gemma-2-27B-It 0.245 0.300 0.190 0.658 0.709 0.607 0.596 0.652 0.540 0.533 0.588 0.478
Qwen2.5-32B-Instruct 0.294 0.337 0.251 0.680 0.722 0.638 0.621 0.674 0.568 0.560 0.604 0.515
Qwen2.5-72B-Instruct 0.281 0.319 0.244 0.685 0.734 0.634 0.621 0.677 0.564 0.569 0.619 0.518
Llama-3.3-70B-Instruct 0.307 0.359 0.255 0.698 0.749 0.647 0.632 0.691 0.574 0.589 0.643 0.536
Gemini-Exp-1206 0.357 0.410 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
GPT-4O-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.650 0.698 0.602
Claude-3-5-Sonnet-20241022 0.444 0.525 0.362 0.727 0.784 0.669 0.692 0.757 0.626 0.652 0.715 0.587
Deepseek-V3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.680
Table 3 provides a comprehensive evaluation of CodeIF. The results are measured using four key metrics: Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). More detailed results are in Appendix 8.2.
Result Overview
DeepSeek-V3 and Claude-3-5-Sonnet-20241022 consistently achieve top scores across all difficulty levels and evaluation metrics, excelling in complex tasks. However, even the highest CSR on the Hard dataset is only 0.362, highlighting room for improvement in handling strict constraints.
Model Scale and Performance Trends
Larger models generally perform better across all metrics. In the Qwen2.5 series (Qwen2.5-1.5B-Instruct to Qwen2.5-72B-Instruct), performance improves with model size. However, this trend is not universal. In the Llama3 series, larger models do not always achieve the best results, suggesting that architecture design, training data quality, and optimization strategies are equally critical factors.
Open-Source vs. Closed-Source Models
Closed-source models like GPT-4O and Claude-3-5-Sonnet-20241022 outperform their open-source counterparts, maintaining higher accuracy under complex constraints. While some large open-source models (e.g., Qwen2.5-72B-Instruct) show competitive results, they still lag behind. This performance gap likely stems from advantages in data quality, optimization techniques, and reinforcement learning from human feedback (RLHF), which proprietary models leverage for enhanced instruction-following precision.
Impact of Task Difficulty
Table 3 shows that as task complexity increases (from Easy to Hard), the performance of nearly all models declines. For example, GPT-4O-2024-11-20 achieves a CSR of 0.441 on Easy tasks, but this drops to 0.325 on Hard tasks. This trend confirms that higher constraint levels impose greater challenges, reducing instruction-following accuracy. These findings align with previous research, which suggests that models’ ability to adhere to instructions deteriorates as task difficulty increases.
5 In-Depth Analysis
Figure 7: Performance of different LLMs on the CodeIF evaluation across instruction categories, measured by SSR.
5.1 Performance Analysis Across Instruction Types
Figure 7 illustrates the performance of LLMs across different instruction categories, revealing notable disparities among models. DeepSeek-V3 consistently achieves the highest scores across most evaluation dimensions, with a particularly strong performance in combination tasks (0.831), highlighting its effectiveness in handling complex instructions. However, its strengths are not uniform; it excels in global structure control but shows relative weaknesses in variable handling, likely reflecting its specific optimization strategies.
Conversely, Meta’s Llama series demonstrates a significant correlation between model size and performance. While smaller configurations (Llama-3.2-1B-Instruct) score lower, larger variants (Llama-3.3-70B-Instruct) perform significantly better, showcasing substantial gains with increased parameter counts. However, model size alone is not the sole determinant of performance. Comparisons between similarly sized models, such as Google’s Gemma and Meta’s Llama, indicate distinct strengths in specific metrics, suggesting that architectural design and training methodologies also play a crucial role.
5.2 Cross-Language Performance Analysis of LLMs
Figure 8: SSR scores of LLMs across different programming languages in the CodeIF evaluation.
Figure 8 compares the performance of leading LLMs across four programming languages: C++, Java, Python, and Go, highlighting key trends at both the model and language levels.
At the model level, DeepSeek-V3 consistently outperforms others, achieving the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), with an RSR of 0.787 in Java, likely benefiting from pretraining optimizations for code continuity. Claude-3-5-Sonnet excels in Java, attaining the highest CSR (0.504) and RSR (0.749), indicating specialized mechanisms for handling complex constraints. However, its SSR in Python (0.703) is relatively lower, suggesting challenges in cross-language generalization. GPT-4O demonstrates balanced performance across languages, ranking second in CSR (0.355) and RSR (0.682) for Python, with minimal variation across different programming paradigms (coefficient of variation CV = 0.18).
At the language level, C++ presents the most significant challenges, likely due to its complex template metaprogramming, which increases constraint difficulty. Java exhibits the highest inter-model variance, with Claude-3-5-Sonnet standing out, possibly due to optimizations tailored to its strong type system. Go achieves the highest overall SSR, benefiting from its simple syntax, but RSR fluctuates significantly, reflecting differences in how models handle concurrency constraints.
These findings highlight variations in LLMs’ ability to generalize across languages and suggest potential areas for optimization, such as improving handling of dependency constraints and enhancing consistency across different programming paradigms.
5.3 Analysis of Instruction Adherence Deviations
Analysis of model-generated responses reveals frequent deviations from specific instructions, particularly in naming conventions and formatting constraints. Models often overlook global formatting rules, such as line limits and character restrictions, when handling general structure instructions.
Moreover, adherence to naming conventions remains inconsistent. For instance, even when explicitly instructed to use PascalCase for variable names, models frequently generate results in lowercase or underscore-separated formats (e.g., transforming current_power into CurrentPower incorrectly).
A particularly notable issue is the tendency to disregard prohibitive instructions. For example, when explicitly instructed to avoid if statements in decision-making—favoring ternary operators or data structures (e.g., dictionaries)—models still generate outputs containing if statements, indicating gaps in constraint enforcement.
5.4 Enhance Model Instruction Compliance
Insights from Appendix Table 5 highlight key strategies for improving instruction adherence in code generation. Supervised Fine-Tuning (SFT) has proven effective, particularly in the Llama series, while model scale remains a crucial factor—larger models like Qwen2.5-72B-Instruct consistently outperform smaller counterparts such as Qwen2.5-3B-Instruct in instruction-following accuracy.
To enhance adherence, several approaches can be employed. First, classifying instructions by priority ensures strict enforcement of hard constraints (e.g., syntax restrictions) while allowing flexibility for soft guidelines (e.g., coding styles). Patterned code generation can provide structured alternatives, such as replacing conditional statements with lookup tables or state machines. Integrating a naming convention engine allows automated conversion of variable and class names (e.g., transforming snake_case to PascalCase). Additionally, leveraging Abstract Syntax Tree (AST) analysis helps detect and transform prohibited structures, such as replacing for loops with while loops or recursive calls. Finally, conflict resolution mechanisms can be introduced to handle contradictory user instructions, such as providing alternative solutions when certain language features are unavailable (e.g., requesting switch-case in Python).
6 Related Work
Code generation and instruction-following are pivotal capacities under examination in AI research (Feng et al., 2020; Sun et al., 2024; Luo et al., 2024; Wang et al., 2023b; Kim et al., 2018; Li et al., 2023; Lu et al., 2021; Li et al., 2022; Wei et al., 2023; Nijkamp et al., 2023b; Zhuo et al., 2024; Jain et al., 2024; Nijkamp et al., 2023a; Zhang et al., 2023; Allal et al., 2023; Lozhkov et al., 2024a; Roziere et al., 2023; Lozhkov et al., 2024b; Wang et al., 2021; Yan et al., 2023). Several benchmarks have been devised to appraise these capabilities in large-scale models. For code generation, benchmarks like McEval (Chai et al., 2024), FullStackBench (Liu et al., 2024), Repocoder (Zhang et al., 2023), Repobench (Liu et al., 2023), and LiveCodeBench (Jain et al., 2024) have been notable. Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench (Qin et al., 2024), CFBench (Zhang et al., 2024), Instruct-following (Zhou et al., 2023), and FollowBench (Jiang et al., 2024), each tailored to assess different aspects of following instructions given to models.
7 Conclusion
This study introduces CodeIF, a benchmark for evaluating the instruction-following capabilities of LLMs in code generation. Covering Java, Python, Go, and C++, CodeIF constructs a diverse test set with constraints ranging from global to specific variables. It introduces novel evaluation metrics—Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR)—to assess multi-constraint handling across multiple dimensions.