### v3

# CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation

Abstract
With the rapid advancement of Large Language Models (LLMs), the demand for robust instruction-following capabilities in code generation tasks has grown significantly. Code generation not only facilitates faster prototyping and automated testing, but also augments developer efficiency through improved maintainability and reusability of code. In this paper, we introduce CodeIF, the first benchmark specifically designed to assess the abilities of LLMs to adhere to task-oriented instructions within diverse code generation scenarios. CodeIF encompasses a broad range of tasks, including function synthesis, error debugging, algorithmic refactoring, and code explanation, thereby providing a comprehensive suite to evaluate model performance across varying complexity levels and programming domains. We conduct extensive experiments with LLMs, analyzing their strengths and limitations in meeting the demands of these tasks. The experimental results offer valuable insights into how well current models align with human instructions, as well as the extent to which they can generate consistent, maintainable, and contextually relevant code. Our findings not only underscore the critical role that instruction-following LLMs can play in modern software development, but also illuminate pathways for future research aimed at enhancing their adaptability, reliability, and overall effectiveness in automated code generation.  [^1] CodeIF data and code are publicly available: https://github.com/lin-rany/codeIF.
CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation
Kaiwen Yan1 *, Hongcheng Guo1 * †, Xuanqing Shi2 Shaosheng Cao3, Donglin Di2, Zhoujun Li1 1Beihang University, 2Tsinghua University, 3Xiaohongshu lin_rany@foxmail.com, hongchengguo@buaa.edu.cn, sxq23@mails.tsinghua.edu.cn caoshaosheng@xiaohongshu.com, donglin.ddl@gmail.com, lizj@buaa.edu.cn
footnotetext: *Equal contributionfootnotetext: Corresponding Author
1 Introduction
With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation. While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge Liu et al. (2024); Zhang et al. (2023). To drive progress in this field, a comprehensive and systematic evaluation framework is essential Jiang et al. (2024); Zhou et al. (2023).
This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation. Built upon insights from existing evaluation sets like McEval Chai et al. (2024) and FullStackBench Liu et al. (2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++. It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
To ensure rigorous assessment, we propose four novel evaluation metrics: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics measure models’ ability to handle multi-constraint problems from different perspectives, including full compliance, average constraint satisfaction, logical coherence, and consistency in instruction execution. By offering a structured evaluation framework, CodeIF provides valuable insights into the current state and future direction of LLM-driven code generation. Overall, our contributions are mainly four-fold:
  1. 1.
    Innovative Benchmark. We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation. CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
  2. 2.
    Automated High-Quality Instruction Generation. Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists. This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
  3. 3.
    Novel Evaluation Metrics. We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks. These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
  4. 4.
    Extensive Evaluation and Analysis. We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels. Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
2 CodeIF
Overview:
Figure 1: The construction process of CodeIF. The first step involves the construction of constraint instructions, followed by the assembly of the dataset, and finally the construction of dependencies between instructions.
As shown in Figure 1, CodeIF is constructed by collecting and refining constraint instructions from real code generation tasks, then combining these tasks with LLM outputs and human review to create a high-quality evaluation dataset.
2.1 Building
The construction of the CodeIF dataset involves two phases: collecting constraint instructions (Section 2.2) and processing data to create the final CodeIF evaluation dataset (Section 2.3).
2.2 Constraint Instructions Collection
The first phase of our work centers on code generation, constructing the CodeIF evaluation dataset through two steps: (1) collecting and verifying constraint instructions, and (2) using them for dataset generation.
We analyze benchmarks like McEval Chai et al. (2024) and FullStackBench Liu et al. (2024) to develop an instruction system spanning eight categories, each targeting specific aspects of code generation for a fine-grained assessment of LLMs’ instruction-following abilities. Constraints are decomposed into atomic instructions with explicit directives, enabling objective binary evaluation (yes/no) and minimizing subjective interpretation. The eight categories cover both architectural-level specifications and variable-level implementation controls, ensuring comprehensive constraint coverage. Specifically, the Global category evaluates adherence to overarching specifications, while Structural Control focuses on control structures (e.g., loops, conditionals) and data structures. Variable constraints assess naming and initialization. Higher abstraction levels include Interface, Function, and Class constraints for program components, while the File category tests cross-file dependencies and external library handling. The Combination category integrates constraints across dimensions, challenging models with complex scenarios.
Figure 2: CodeIF Constraints Instruction Distribution
Figure 2 presents CodeIF’s distribution across programming languages and categories. The evaluation system features 8 categories and 50 fine-grained constraint instructions, systematically assessing LLMs’ code generation performance. By organizing constraints clearly, the system identifies strengths and weaknesses, guides optimization, and advances automated code generation. The full list of constraints is in Appendix 4.
2.3 Data Construction
Multi-Language and Difficulty-Differentiated Benchmark Design
To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval Chai et al. (2024) and FullStackBench Liu et al. (2024). These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
To further refine the evaluation, we categorize tasks into two difficulty levels: Easy and Hard. The Hard set includes longer, more intricate instruction lists, designed to rigorously test LLMs’ ability to handle complex constraints.
Automated Generation of Constraint Instructions
We used large language models (LLMs) like GPT-4 to create task-specific instruction lists for code generation tasks. We prepared 20 detailed examples and formulated concise atomic instructions for accuracy. These examples guided LLMs in refining tasks and streamlining instructions to enhance clarity and output quality.
Constructing Instruction Dependencies
We utilized LLMs to map dependencies between atomic constraints, improving our evaluation framework’s precision and verification accuracy. By understanding the dependencies among instructions, we outlined clear steps for tasks like function creation, which involve naming the function, defining parameter types, and coding the body. Incorporating these dependencies enhances our evaluation system, more accurately assessing the model’s capability with complex instructions and identifying areas for improvement. Figure 3 illustrates a CodeIF task with its instruction sequence and dependencies.
Figure 3: Specific cases of the CodeIF dataset, ’Task’ denotes the specific generation task, ’Type’ refers to the type of constraint, and ’Dependence’ indicates the prerequisite constraints for this constraint.
2.4 Data Analysis
Set Num Avg.Instr Go Python Java C++
Easy 600 11.99 127 165 176 132
Hard 600 13.80 103 183 177 137
Full 1200 12.90 230 348 353 269
Table 1: CodeIF dataset statistics, showing the statistical information of different difficulty classifications. Avg.Instr represents the average length of the atomic constraint instruction list.
CodeIF Static Analysis
Figure 4: The distribution of atomic instruction list lengths in datasets of different difficulties.
Table 1 categorizes the dataset into three difficulty levels: Easy, Hard, and Full. Both Easy and Hard sets contain 600 tasks, while the Full dataset combines them, totaling 1,200 tasks across Go, Python, Java, and C++. Java has the most tasks (353), followed by Python (348), C++ (269), and Go (230). The Easy set averages 11.99 instructions per task, the Hard set 13.8, and the Full dataset 12.9, reflecting increasing complexity. Figure 4 shows task length distribution.
Set Global Structural Control Variable Interface Function Class File Combination
Easy 1638 1008 1336 427 569 544 723 953
Hard 1890 1193 1479 505 659 623 802 1142
Full 3528 2201 2815 932 1228 1167 1525 2095
Table 2: CodeIF dataset statistics information, showing the distribution of atomic restriction instruction categories under different difficulty classifications.
Constraint Instruction Analysis
Table 2compares instruction distribution across difficulty levels. The Hard set consistently has more instructions per category than the Easy set, with the Global category averaging 2.5 instructions in Easy and over 3 in Hard. This indicates greater challenges for models as task complexity rises. More analysis is in Appendix D.
3 Metrics
Ensuring that large language models (LLMs) accurately follow instructions is crucial for code generation. To precisely evaluate this capability, we introduce four novel metrics designed to assess how LLMs handle code generation tasks with multiple constraints: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics provide a comprehensive evaluation from different perspectives.
For a dataset with $m$ problems, each problem contains a set of $n_{i}$ constraints. We define CSR and SSR as follows:
Completely Satisfaction Rate (CSR):
$\text{CSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\prod_{j=1}^{n_{i}}r_{i,j}\right)$ (1)
where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied.
Soft Satisfaction Rate (SSR):
$\text{SSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}r_{i,j}}{n_{i}}\right)$ (2)
SSR evaluates the average proportion of constraints satisfied per problem, providing a more flexible assessment.
Rigorous Satisfaction Rate (RSR)
In code generation, some constraints depend on prior instructions, particularly in Combination constraints. To account for dependencies, we define RSR as:
$\text{RSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}\left[r_{i,j}\cdot\prod_{k\in D_{i,j}}r_{i,k}\right]}{n_{i}}\right)$ (3)
where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on.
Consistent Continuity Satisfaction Rate (CCSR)
In many code generation tasks, maintaining continuous adherence to instructions is essential. To measure this ability, we define CCSR as:
$\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$.
4 Experiment
4.1 Experimental Setup
The temperature coefficient is set to 0 to ensure output determinism, with a maximum generation length of 4096 tokens. All other settings follow the official default parameters for each model. Commercial API models are accessed through the latest available interface as of December 2024. All experiments are conducted using the official API and 8 H800(80G).
4.2 Automatic Evaluation
To ensure robust evaluation, we used LLMs and human experts to verify model adherence to atomic constraints. Constraints were decomposed into atomic elements, enabling objective binary evaluations (Yes/No) over subjective judgments. Following FairEval Wang et al. (2023a), GPT-4-1106-Preview was the primary evaluation tool (prompt details in Appendix A). Three domain experts manually annotated 150 stratified samples. Statistical analysis showed strong agreement, with Pearson correlations of 0.87 (LLM-human) and 0.85 (inter-human), confirming high consistency. Baselines are in Appendix C.
4.3 Main Experiments
Models CSR SSR RSR CCSR
Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
Llama-3.2-1B-Instruct 0.034 0.046 0.022 0.218 0.277 0.159 0.182 0.231 0.132 0.152 0.197 0.107
Llama-3.1-8B-Instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.340 0.370 0.444 0.295
Qwen2.5-Coder-7B-Instruct 0.142 0.198 0.087 0.514 0.590 0.438 0.453 0.533 0.373 0.390 0.463 0.318
Qwen2.5-7B-Instruct 0.153 0.201 0.104 0.535 0.599 0.471 0.475 0.546 0.405 0.416 0.479 0.353
Ministral-8B 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.490 0.371
Gemma-2-9B-It 0.171 0.210 0.131 0.573 0.642 0.504 0.513 0.587 0.440 0.445 0.508 0.383
Qwen2.5-Coder-32B-Instruct 0.365 0.422 0.307 0.736 0.767 0.704 0.679 0.723 0.635 0.634 0.669 0.599
Gemma-2-27B-It 0.245 0.300 0.190 0.658 0.709 0.607 0.596 0.652 0.540 0.533 0.588 0.478
Qwen2.5-32B-Instruct 0.294 0.337 0.251 0.680 0.722 0.638 0.621 0.674 0.568 0.560 0.604 0.515
Qwen2.5-72B-Instruct 0.281 0.319 0.244 0.685 0.734 0.634 0.621 0.677 0.564 0.569 0.619 0.518
Llama-3.3-70B-Instruct 0.307 0.359 0.255 0.698 0.749 0.647 0.632 0.691 0.574 0.589 0.643 0.536
Gemini-Exp-1206 0.357 0.410 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
GPT-4O-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.650 0.698 0.602
Claude-3-5-Sonnet-20241022 0.444 0.525 0.362 0.727 0.784 0.669 0.692 0.757 0.626 0.652 0.715 0.587
Deepseek-V3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.680
Table 3: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
Table 3 evaluates CodeIF using four metrics: CSR, SSR, RSR, and CCSR. Detailed results are in Appendix B.
Overview.
DeepSeek-V3 and Claude-3-5-Sonnet-20241022 lead across metrics, excelling in complex tasks. However, the highest CSR on Hard tasks is just 0.362, showing challenges in meeting strict constraints.
Model Scale Trends.
Larger models generally perform better, as seen in Qwen2.5 series. However, the Llama3 series shows inconsistent results, highlighting the importance of architecture, data quality, and optimization.
Open vs. Closed Models.
Closed-source models like GPT-4O and Claude-3-5 outperform open-source models, especially under complex constraints. While large open-source models (e.g., Qwen2.5-72B-Instruct) are competitive, they lag due to differences in data quality and RLHF techniques.
Task Difficulty Impact.
Performance drops with increasing task complexity. For instance, GPT-4O’s CSR falls from 0.441 on Easy tasks to 0.325 on Hard tasks, highlighting the challenge of strict constraints.
5 In-Depth Analysis
Figure 5: Performance of different LLMs on the CodeIF evaluation across instruction categories, measured by SSR.
5.1 Performance Analysis Across Instruction Types
Figure 5 compares LLM performance across instruction categories, revealing notable variations. DeepSeek-V3 leads overall, excelling in combination tasks (0.831) and global structure control, though weaker in variable handling, reflecting its optimization focus. Meta’s Llama series shows a clear correlation between model size and performance, with larger variants (Llama-3.3-70B-Instruct) outperforming smaller ones (Llama-3.2-1B-Instruct). However, size alone is not decisive; comparisons with similarly sized models like Google’s Gemma highlight the role of architecture and training methods in shaping performance.
5.2 Cross-Language Performance Analysis of LLMs
Figure 6: SSR scores of LLMs across different programming languages in the CodeIF evaluation.
Figure 6 compares the performance of leading LLMs across C++, Java, Python, and Go, highlighting trends at both model and language levels. At the model level, DeepSeek-V3 leads with the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), and an RSR of 0.787 in Java. Claude-3-5-Sonnet excels in Java with the highest CSR (0.504) and RSR (0.749), but shows lower SSR in Python (0.703). GPT-4O demonstrates balanced performance, ranking second in Python’s CSR (0.355) and RSR (0.682), with minimal variance (CV = 0.18). At the language level, C++ is the most challenging due to complex template metaprogramming. Java shows high inter-model variance, with Claude-3-5-Sonnet performing best. Go achieves the highest SSR but fluctuates in RSR. These results highlight cross-language generalization differences and suggest optimization areas like dependency handling and paradigm consistency.
5.3 Analysis of Instruction Adherence Deviations
Analysis of model-generated responses shows frequent deviations from instructions, especially in naming conventions and formatting constraints. Models often ignore global formatting rules, such as line limits, and inconsistently follow naming conventions. For example, when instructed to use PascalCase, models sometimes output lowercase or underscore-separated formats (e.g., incorrectly transforming current_power into CurrentPower). A notable issue is the disregard for prohibitive instructions. For instance, models often use if statements despite being instructed to avoid them in favor of ternary operators or data structures like dictionaries, revealing gaps in constraint enforcement.
5.4 Improving Instruction Compliance
Appendix Table 5 highlights strategies to improve adherence. Supervised Fine-Tuning (SFT) proves effective, especially in the Llama series, while larger models like Qwen2.5-72B-Instruct outperform smaller ones in instruction-following accuracy. Key improvements include prioritizing hard constraints (e.g., syntax rules) over soft guidelines (e.g., coding styles). Patterned code generation can replace conditional statements with lookup tables or state machines. A naming convention engine can automate variable name formatting (e.g., converting snake_case to PascalCase). Abstract Syntax Tree (AST) analysis can detect and transform prohibited structures, such as replacing for loops with while loops. Conflict resolution mechanisms can address contradictory instructions, offering alternative solutions when certain language features are unavailable (e.g., using Python’s alternatives to switch-case).
6 Related Work
Code generation and instruction-following are pivotal capacities under examination in AI research Feng et al. (2020); Sun et al. (2024); Luo et al. (2024); Wang et al. (2023b); Kim et al. (2018); Li et al. (2023); Lu et al. (2021); Li et al. (2022); Wei et al. (2023); Nijkamp et al. (2023b); Zhuo et al. (2024); Jain et al. (2024); Nijkamp et al. (2023a); Zhang et al. (2023); Allal et al. (2023); Lozhkov et al. (2024a); Roziere et al. (2023); Lozhkov et al. (2024b); Wang et al. (2021); Yan et al. (2023). Several benchmarks have been devised to appraise these capabilities in large-scale models. For code generation, benchmarks like McEval Chai et al. (2024), FullStackBench Liu et al. (2024), Repocoder Zhang et al. (2023), Repobench Liu et al. (2023), and LiveCodeBench Jain et al. (2024) have been notable. Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench Qin et al. (2024), CFBench Zhang et al. (2024), Instruct-following Zhou et al. (2023), and FollowBench Jiang et al. (2024), each tailored to assess different aspects of following instructions given to models.
7 Conclusion
This study introduces CodeIF, a benchmark for evaluating the instruction-following capabilities of LLMs in code generation. Covering Java, Python, Go, and C++, CodeIF constructs a diverse test set with constraints ranging from global to specific variables. It introduces novel evaluation metrics—Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR)—to assess multi-constraint handling across multiple dimensions.
8 Limitations
Limited Language Support. CodeIF includes key languages like Java, Python, Go, and C++, but excludes popular ones like JavaScript, Ruby, and Swift. Expanding language coverage would improve its applicability in diverse contexts.
Static Evaluation Focus. CodeIF focuses mainly on static code properties, such as structure and naming conventions, while overlooking dynamic factors like runtime behavior, performance, and debugging. Including dynamic evaluation would better reflect real-world development challenges.
Uniform Metric Weighting. The metrics (CSR, SSR, RSR, CCSR) treat all constraints equally, which may not align with practical priorities. For example, syntactic correctness is often more critical than naming conventions. Introducing weighted scoring could enhance the interpretability of model performance.
Appendix A Prompt Template
Prompt for Instruction Generation You are an instruction compliance evaluator, required to assess the instruction compliance ability of large models. Therefore, you need to generate a series of data for the code generation instruction detection of large models. [Input Format]
I will input a series of data, and you need to generate a dictionary based on these data, which includes two fields “question” and “instruction_list”
Original question: {Original question} Original instruction list:
{instruction_list}
Input Explanation
The original question is the original question. It contains the original code generation problem.
The original instruction list is the original instruction list. It contains randomly generated code compliance instructions. Some instructions will contain directive keywords that need to be replaced and are wrapped in {{}}. Return Format
Return a json data, do not have extra output. The returned dictionary contains two fields: “question” and “instruction_list”
The format is as follows:
    "question": "Optimized question",
    "instruction_list": [
            "instruction_id": "id1",
            "instruction": "Instruction 1"
Explanation
“question”: It is the optimized question, which does not contain any directive instructions, only contains the explanation of the original question. It does not contain any restrictions on the code. Move the instructions in the question to the instruction list
“instruction_list”: It is the optimized instruction list. You should optimize according to the meaning of the question. More in line with the meaning of the question. Instead of directly outputting the original instruction list, note that you should replace all directive keywords that need to be replaced and are wrapped in , and the final output should not contain directive keywords that need to be replaced. Generation Requirements
question: Please generate the optimized question based on the following data, which does not contain any directive instructions, only contains the core content of the original question.
instruction_list: Originated from the input original instruction list. If there are instructions that completely conflict with the meaning of the question or instructions that conflict with each other You should delete as little as possible, you should modify more. Please replace according to the content in the original instruction_list, you should delete as little as possible. Unless it is contradictory instructions, or instructions that cannot be achieved at all, if you only need to generate additional code to meet the requirements, you can replace it.
Prompt for Code Generation As a programming assistant, your task is to generate code snippets based on the user question and instructions given below: Please consider the following points while generating the code snippet: - Make sure you follow the user instructions word to word. If the instruction says to use a specific language or a specific method, use exactly that. - Your output should be a valid code snippet in the programming language indicated in the question or the instructions. - Pay close attention to the syntax and formatting rules of the programming language that you are using. The code should be well-formatted and easy to read. - Make sure that the code snippet you are generating is efficient and is not overly complicated. Output Format: The output should be a valid, well-formatted, and efficient code snippet that adheres to the above question and instructions. Task information User Question: {question} Instructions: {instructions_str} Please generate the code snippet based on the above information:
Prompt for Answer Judgment As a programming assistant, your task is to evaluate whether the generated code strictly follows the instructions given in light of the user’s problem and directives. You need to return a list of the same length as the instructions, containing only ’Yes’ and ’No’, indicating whether the model adhered to each specific instruction. Consider the following when making judgments: - You must strictly follow the user’s instructions. If the instruction requires the use of a specific language or method, you must explicitly check if the code utilizes it. - Your output should be a list of the same length as the instructions, containing only ’Yes’ and ’No’. - Pay close attention to the programming language syntax and formatting rules you are evaluating. The code should be neatly organized and easy to read. - The list you generate should be valid and not overly complex. Task Information User question:
{question}
Instructions:
{instructions_str}
Model-generated response:
{generated_code}
Based on the information provided, determine whether the model has followed the instructions, and return a list of the same length as the instructions, containing only ‘Yes’ and ‘No’. Please note!!! Your output should only contain the list, with no other content. The items in the list should only be ‘Yes’ and ‘No’, with no other words included.
Appendix B More Resluts
ID Type Instruction Format Format Keys
1 global Your entire response should be written in {programming_language}, the use of other programming languages is not allowed. ["programming_language"]
2 global Your code lines should not exceed {characters_num} characters. ["characters_num"]
3 global Your code should use global variables. []
4 global Your code should not use global variables. []
5 global Your function should have at most {parameter_count} parameters. ["parameter_count"]
6 global Your code should not have more than {function_count} functions. ["function_count"]
7 global Your code should not have more than {class_count} classes. ["class_count"]
8 global Your code should not use the {keyword} keyword. ["keyword"]
9 global Your function should not exceed {line_num} lines. ["line_num"]
10 global Your answer in total should not exceed {line_num} lines. ["line_num"]
11 global Your code should use the {keyword} keyword. ["keyword"]
12 structural control Your code should use data structure {data_structure}. ["data_structure"]
13 structural control Your code should not use data structure {data_structure}. ["data_structure"]
14 structural control Your code should use for-loop. []
15 structural control Your code should not use for-loop. []
16 structural control Your code should use while-loop. []
17 structural control Your code should not use while-loop. []
18 structural control Your code should use if statement for decision making. []
19 structural control Your code should not use if statement for decision making. []
20 structural control Your code should use switch statement for decision making. []
21 structural control Your code should not use switch statement for decision making. []
22 variable Your code should define a variable named {variable_name}. ["variable_name"]
23 variable Your code should define an enumeration named {enumeration_name} ["enumeration_name"]
24 variable The variable names in your code should follow the {naming_convention} naming convention ["naming_convention"]
25 variable Variable {variable_name}, type should be {variable_type}. ["variable_name", "variable_type"]
26 variable Variable {variable_name}, should be a global variable. ["variable_name"]
27 variable Variable {variable_name}, should not be a global variable. ["variable_name"]
28 variable Variable {variable_name}, the initial value should be {variable_value}. ["variable_name", "variable_value"]
29 variable Variable {variable_name}, should be a constant. ["variable_name"]
30 variable Variable {variable_name} should not be a constant. ["variable_name"]
31 function Your code should include a function named {function_name}. ["function_name"]
32 function The function names in your code should follow the {naming_convention}. naming convention ["naming_convention"]
33 function Your code should not use any functions from the {disallowed_function_list}. ["disallowed_function_list"]
34 interface Your code should define an interface named {interface_name}. ["interface_name"]
35 interface The interface names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
36 class Your code should define a class named {class_name}. ["class_name"]
37 class The class names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
38 file Your code should be organized in a package named {package_name}. ["package_name"]
39 file Your code should import the following libraries {library_list}. ["library_list"]
40 file Your code should use the function {function_name} from the library {library_name}. ["function_name", "library_name"]
41 file Your code should not use the following libraries {disallowed_library_list}. ["disallowed_library_list"]
42 combination You should initialize an object named {object_name} as an instance of the {class_name} class using {parameters_name_list} for initialization. ["object_name", "class_name", "parameters_name_list"]
43 combination You should define an interface named {interface_name} that includes these methods {method_name_list}. ["interface_name", "method_name_list"]
44 combination Your code should define a class named {class_name} that implements the {interface_name} interface. ["class_name", "interface_name"]
45 combination In your code, the class {class_name} should have these properties {properties_name_list}. ["class_name", "properties_name_list"]
46 combination In your code, the class {class_name} should have these methods {method_name_list}. ["class_name", "method_name_list"]
47 combination The function {function_name} should take {parameter_name_list} as parameters. ["function_name", "parameter_name_list"]
48 combination The function {function_name} should return a {return_type} as its result. ["function_name", "return_type"]
49 combination Your code should be organized in a package named {package_name}, which should contain these classes {class_name_list}. ["package_name", "class_name_list"]
50 combination Your code should be organized in a package named {package_name}, which should contain these functions {function_name_list}. ["package_name", "function_name_list"]
Table 4: Constraint Instruction Table
Models CSR SSR RSR CCSR
Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
Llama-3.2-1b-instruct 0.034 0.046 0.022 0.218 0.277 0.159 0.182 0.231 0.132 0.152 0.197 0.107
Qwen2.5-1.5b-instruct 0.034 0.053 0.015 0.265 0.334 0.197 0.222 0.282 0.162 0.181 0.234 0.128
Qwen2.5-coder-1.5b-instruct 0.058 0.086 0.03 0.358 0.436 0.281 0.301 0.371 0.233 0.251 0.314 0.189
Qwen2.5-3b-instruct 0.078 0.109 0.046 0.415 0.489 0.34 0.357 0.432 0.282 0.299 0.364 0.233
Llama-3.2-3b-instruct 0.101 0.137 0.065 0.396 0.473 0.318 0.344 0.419 0.268 0.305 0.375 0.235
GPT-3.5-turbo 0.102 0.14 0.065 0.41 0.467 0.353 0.362 0.42 0.303 0.314 0.369 0.259
Qwen2.5-coder-3b-instruct 0.097 0.142 0.051 0.445 0.529 0.359 0.383 0.464 0.301 0.33 0.401 0.258
Llama-3.1-8b 0.129 0.178 0.08 0.452 0.551 0.353 0.402 0.497 0.306 0.352 0.44 0.263
Llama-3.1-8b-instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.34 0.37 0.444 0.295
Qwen2.5-coder-7b-instruct 0.142 0.198 0.087 0.514 0.59 0.438 0.453 0.533 0.373 0.39 0.463 0.318
Ministral-3b 0.127 0.162 0.092 0.526 0.591 0.46 0.458 0.527 0.39 0.4 0.458 0.342
Phi-3.5-mini-128k-instruct 0.154 0.217 0.09 0.514 0.635 0.391 0.456 0.574 0.337 0.405 0.51 0.299
Qwen2.5-7b-instruct 0.153 0.201 0.104 0.535 0.599 0.471 0.475 0.546 0.405 0.416 0.479 0.353
Ministral-8b 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.49 0.371
Gemma-2-9b-it 0.171 0.21 0.131 0.573 0.642 0.504 0.513 0.587 0.44 0.445 0.508 0.383
Llama-3.1-70b 0.196 0.232 0.16 0.61 0.664 0.555 0.545 0.607 0.482 0.482 0.533 0.43
Qwen2.5-coder-14b-instruct 0.218 0.276 0.16 0.596 0.667 0.525 0.539 0.614 0.463 0.483 0.55 0.416
Qwen2.5-14b-instruct 0.238 0.279 0.198 0.61 0.676 0.543 0.557 0.628 0.486 0.498 0.565 0.431
Gemini-2.0-flash-exp 0.254 0.29 0.218 0.615 0.648 0.583 0.556 0.593 0.518 0.514 0.547 0.481
Gemma-2-27b-it 0.245 0.3 0.19 0.658 0.709 0.607 0.596 0.652 0.54 0.533 0.588 0.478
Llama-3.1-70b-instruct 0.265 0.3 0.229 0.675 0.723 0.627 0.612 0.667 0.556 0.559 0.601 0.516
Qwen2.5-32b-instruct 0.294 0.337 0.251 0.68 0.722 0.638 0.621 0.674 0.568 0.56 0.604 0.515
Qwen2.5-72b-instruct 0.281 0.319 0.244 0.685 0.734 0.634 0.621 0.677 0.564 0.569 0.619 0.518
Codestral-2501 0.28 0.339 0.219 0.683 0.748 0.617 0.621 0.691 0.551 0.571 0.633 0.507
Phi-4 0.312 0.361 0.262 0.698 0.735 0.66 0.635 0.681 0.589 0.589 0.631 0.546
Llama-3.3-70b-instruct 0.307 0.359 0.255 0.698 0.749 0.647 0.632 0.691 0.574 0.589 0.643 0.536
GPT-4o-mini 0.292 0.348 0.237 0.731 0.78 0.682 0.665 0.724 0.606 0.609 0.66 0.559
GPT-4o 0.338 0.392 0.283 0.721 0.77 0.671 0.665 0.721 0.609 0.616 0.668 0.563
Qwen2.5-coder-32b-instruct 0.365 0.422 0.307 0.736 0.767 0.704 0.679 0.723 0.635 0.634 0.669 0.599
Gemini-exp-1206 0.357 0.41 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
Gemini-1.5-pro 0.351 0.383 0.318 0.763 0.794 0.732 0.704 0.744 0.663 0.647 0.679 0.615
GPT-4o-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.65 0.698 0.602
Claude-3-5-sonnet-20241022 0.444 0.525 0.362 0.727 0.784 0.669 0.692 0.757 0.626 0.652 0.715 0.587
Deepseek-coder 0.41 0.45 0.37 0.805 0.836 0.773 0.749 0.791 0.707 0.699 0.732 0.666
Deepseek-v3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.68
Table 5: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
Models Global Structural Control Variable Interface Function Class File Combination
Llama-3.2-1b-instruct 0.186 0.190 0.206 0.144 0.284 0.260 0.198 0.172
Qwen2.5-1.5b-instruct 0.244 0.236 0.221 0.213 0.355 0.315 0.230 0.213
Qwen2.5-coder-1.5b-instruct 0.328 0.304 0.326 0.293 0.436 0.426 0.351 0.304
Qwen2.5-3b-instruct 0.383 0.346 0.412 0.383 0.468 0.481 0.383 0.366
Llama-3.2-3b-instruct 0.344 0.332 0.393 0.376 0.454 0.447 0.363 0.367
GPT-3.5-turbo 0.388 0.344 0.417 0.375 0.467 0.449 0.378 0.352
Qwen2.5-coder-3b-instruct 0.397 0.367 0.438 0.419 0.511 0.507 0.415 0.403
Llama-3.1-8b 0.410 0.355 0.451 0.424 0.500 0.503 0.413 0.413
Llama-3.1-8b-instruct 0.422 0.373 0.482 0.455 0.524 0.499 0.407 0.437
Qwen2.5-coder-7b-instruct 0.479 0.419 0.497 0.502 0.576 0.571 0.492 0.487
Ministral-3b 0.472 0.403 0.527 0.512 0.618 0.609 0.524 0.535
Phi-3.5-mini-128k-instruct 0.461 0.410 0.512 0.531 0.562 0.574 0.485 0.491
Qwen2.5-7b-instruct 0.484 0.425 0.532 0.548 0.616 0.591 0.520 0.520
Ministral-8b 0.497 0.433 0.541 0.570 0.622 0.631 0.527 0.557
Gemma-2-9b-it 0.541 0.498 0.599 0.510 0.659 0.618 0.543 0.511
Llama-3.1-70b 0.558 0.500 0.652 0.653 0.685 0.671 0.545 0.597
Qwen2.5-coder-14b-instruct 0.541 0.467 0.623 0.669 0.645 0.652 0.547 0.594
Qwen2.5-14b-instruct 0.569 0.526 0.652 0.592 0.649 0.644 0.533 0.559
Gemini-2.0-flash-exp 0.555 0.526 0.653 0.666 0.685 0.669 0.564 0.615
Gemma-2-27b-it 0.621 0.569 0.699 0.640 0.722 0.710 0.607 0.637
Llama-3.1-70b-instruct 0.606 0.546 0.722 0.718 0.744 0.738 0.603 0.680
Qwen2.5-32b-instruct 0.637 0.581 0.713 0.712 0.732 0.742 0.601 0.653
Qwen2.5-72b-instruct 0.633 0.570 0.734 0.711 0.727 0.726 0.645 0.686
Codestral-2501 0.617 0.552 0.723 0.718 0.733 0.746 0.651 0.694
Phi-4 0.633 0.586 0.734 0.739 0.721 0.752 0.677 0.710
Llama-3.3-70b-instruct 0.621 0.634 0.733 0.730 0.759 0.738 0.645 0.695
GPT-4o-mini 0.671 0.663 0.787 0.774 0.784 0.783 0.657 0.710
GPT-4o 0.665 0.651 0.742 0.759 0.743 0.759 0.666 0.716
Qwen2.5-coder-32b-instruct 0.683 0.654 0.776 0.763 0.772 0.758 0.695 0.736
Gemini-exp-1206 0.690 0.677 0.780 0.789 0.798 0.809 0.675 0.727
Gemini-1.5-pro 0.718 0.696 0.814 0.800 0.812 0.815 0.672 0.749
GPT-4o-2024-11-20 0.685 0.666 0.784 0.786 0.779 0.785 0.706 0.755
Claude-3-5-sonnet-20241022 0.677 0.678 0.750 0.736 0.742 0.730 0.640 0.692
Deepseek-coder 0.759 0.714 0.850 0.856 0.846 0.847 0.754 0.813
Deepseek-v3 0.780 0.732 0.866 0.876 0.866 0.873 0.762 0.831
Table 6: The performance of various models on CodeIF for different types of instructions
Models CPP Java Python Go
CCS CS SS RS CCS CS SS RS CCS CS SS RS CCS CS SS RS
Llama-3.2-1b-instruct 0.123 0.023 0.185 0.150 0.190 0.037 0.265 0.221 0.179 0.047 0.262 0.223 0.086 0.022 0.117 0.096
Qwen2.5-1.5b-instruct 0.171 0.023 0.250 0.206 0.191 0.026 0.277 0.228 0.197 0.047 0.298 0.257 0.151 0.040 0.216 0.179
Qwen2.5-coder-1.5b-instruct 0.253 0.068 0.348 0.297 0.259 0.055 0.375 0.308 0.263 0.060 0.380 0.328 0.218 0.049 0.309 0.255
Qwen2.5-3b-instruct 0.251 0.046 0.367 0.302 0.310 0.078 0.419 0.367 0.306 0.092 0.435 0.384 0.327 0.093 0.433 0.365
Llama-3.2-3b-instruct 0.284 0.073 0.377 0.313 0.345 0.121 0.435 0.380 0.321 0.112 0.429 0.383 0.244 0.084 0.304 0.265
GPT-3.5-turbo 0.301 0.085 0.388 0.332 0.367 0.134 0.461 0.409 0.265 0.092 0.371 0.334 0.318 0.088 0.412 0.363
Qwen2.5-coder-3b-instruct 0.339 0.103 0.444 0.380 0.338 0.101 0.453 0.391 0.320 0.091 0.446 0.390 0.323 0.093 0.431 0.363
Llama-3.1-8b 0.319 0.115 0.409 0.354 0.366 0.130 0.477 0.420 0.376 0.152 0.485 0.446 0.330 0.110 0.413 0.363
Llama-3.1-8b-instruct 0.328 0.112 0.432 0.375 0.408 0.173 0.503 0.447 0.393 0.147 0.496 0.455 0.325 0.133 0.406 0.365
Qwen2.5-coder-7b-instruct 0.389 0.147 0.505 0.434 0.375 0.118 0.503 0.444 0.400 0.155 0.531 0.475 0.401 0.154 0.516 0.456
Ministral-3b 0.356 0.107 0.473 0.401 0.410 0.150 0.542 0.476 0.404 0.112 0.538 0.481 0.430 0.138 0.544 0.464
Phi-3.5-mini-128k-instruct 0.354 0.108 0.461 0.388 0.426 0.179 0.532 0.478 0.440 0.180 0.559 0.510 0.380 0.131 0.482 0.422
Qwen2.5-7b-instruct 0.401 0.162 0.514 0.448 0.439 0.152 0.559 0.495 0.397 0.147 0.523 0.471 0.429 0.154 0.541 0.485
Ministral-8b 0.400 0.143 0.518 0.439 0.434 0.158 0.560 0.495 0.410 0.142 0.538 0.481 0.494 0.214 0.599 0.532
Gemma-2-9b-it 0.446 0.200 0.560 0.499 0.446 0.164 0.576 0.518 0.380 0.131 0.510 0.456 0.542 0.204 0.678 0.609
Llama-3.1-70b 0.487 0.201 0.598 0.518 0.507 0.232 0.632 0.572 0.425 0.136 0.579 0.521 0.522 0.226 0.635 0.571
Qwen2.5-coder-14b-instruct 0.464 0.224 0.572 0.514 0.478 0.206 0.592 0.535 0.522 0.216 0.653 0.594 0.454 0.235 0.544 0.490
Qwen2.5-14b-instruct 0.481 0.230 0.590 0.533 0.528 0.265 0.639 0.581 0.472 0.188 0.599 0.550 0.511 0.281 0.603 0.557
Gemini-2.0-flash-exp 0.491 0.259 0.587 0.519 0.575 0.309 0.664 0.604 0.468 0.207 0.584 0.533 0.514 0.233 0.619 0.558
Gemma-2-27b-it 0.529 0.271 0.645 0.579 0.551 0.261 0.676 0.616 0.465 0.179 0.604 0.543 0.611 0.289 0.727 0.665
Llama-3.1-70b-instruct 0.535 0.267 0.653 0.578 0.581 0.276 0.685 0.620 0.555 0.251 0.696 0.639 0.557 0.264 0.655 0.596
Qwen2.5-32b-instruct 0.551 0.314 0.655 0.602 0.589 0.330 0.706 0.638 0.522 0.231 0.665 0.609 0.580 0.311 0.690 0.634
Qwen2.5-72b-instruct 0.543 0.297 0.638 0.574 0.580 0.288 0.701 0.633 0.574 0.284 0.702 0.651 0.573 0.249 0.687 0.610
Codestral-2501 0.562 0.307 0.658 0.595 0.583 0.301 0.694 0.632 0.566 0.249 0.693 0.637 0.569 0.261 0.681 0.611
Phi-4 0.570 0.331 0.663 0.601 0.612 0.328 0.719 0.650 0.587 0.295 0.714 0.660 0.577 0.292 0.679 0.613
Llama-3.3-70b-instruct 0.558 0.300 0.652 0.582 0.621 0.348 0.713 0.644 0.572 0.264 0.709 0.653 0.602 0.317 0.712 0.640
GPT-4o-mini 0.582 0.292 0.684 0.615 0.620 0.299 0.738 0.667 0.586 0.261 0.731 0.674 0.661 0.330 0.775 0.707
GPT-4o 0.600 0.337 0.698 0.639 0.652 0.368 0.748 0.693 0.600 0.312 0.723 0.676 0.603 0.332 0.701 0.636
Qwen2.5-coder-32b-instruct 0.633 0.384 0.717 0.658 0.654 0.401 0.753 0.699 0.621 0.342 0.736 0.688 0.624 0.322 0.731 0.661
Gemini-exp-1206 0.640 0.424 0.726 0.672 0.650 0.360 0.755 0.689 0.590 0.290 0.724 0.674 0.677 0.373 0.777 0.710
Gemini-1.5-pro 0.635 0.370 0.741 0.676 0.674 0.379 0.783 0.720 0.610 0.278 0.758 0.706 0.674 0.395 0.764 0.707
GPT-4o-2024-11-20 0.653 0.374 0.741 0.669 0.683 0.434 0.776 0.716 0.612 0.355 0.724 0.682 0.653 0.358 0.747 0.683
Claude-3-5-sonnet-20241022 0.615 0.425 0.684 0.643 0.720 0.504 0.789 0.749 0.611 0.396 0.703 0.674 0.650 0.444 0.716 0.686
Deepseek-coder 0.709 0.441 0.802 0.735 0.731 0.463 0.819 0.764 0.657 0.336 0.791 0.747 0.702 0.403 0.805 0.744
Deepseek-v3 0.725 0.435 0.831 0.762 0.753 0.497 0.839 0.787 0.651 0.315 0.793 0.744 0.722 0.404 0.822 0.76
Table 7: the evaluation results of different languages on CODEIF. The metrics include Consistent Continuity Satisfaction Rate (CCSR), Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), and Rigorous Satisfaction Rate (RSR).
Appendix C Baselines
We evaluate over 30 language models spanning both open-source architectures and commercial APIs. The Meta Llama 3 Series Touvron et al. (2023) contains Llama-3.2-1B/3B/8B/70B-Instruct variants and Llama-3.3-70B-Instruct. Qwen2.5 Series Yang et al. (2024) encompasses Qwen2.5-1.5B/3B/7B/14B/32B/72B-Instruct with dedicated code generation models Qwen2.5-Coder-1.5B/3B/7B/14B/32B-Instruct Hui et al. (2024). Mistral Series Jiang et al. (2023) includes Mistral-3B, Mistral-8B, and the code-specialized Codestral-2501.
The evaluation covers Microsoft’s Phi-3.5-Mini-128K-Instruct (3.8B) and Phi-4 Abdin et al. (2024), along with Google’s Gemma-2-9B/27B-It Team (2024b). DeepSeek Series incorporates DeepSeek-Coder Guo et al. (2024) and DeepSeek-V3 DeepSeek-AI (2024). Commercial APIs include OpenAI’s GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and GPT-4O-2024-11-20 Achiam et al. (2023); Google’s Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro Team (2024a); plus Anthropic’s Claude-3.5-Sonnet-20241022.
Appendix D More Data Analysis
Figure 7: Distribution of atomic instruction list lengths across difficulty levels.
Figure 7 shows the proportion of each instruction category. Global constraints dominate (22.77%), followed by Variable constraints (18.17%). This distribution reflects CodeIF’s balanced focus on high-level structural coherence and fine-grained variable precision, ensuring comprehensive evaluation of code generation capabilities. Figure 8 compares instruction distribution across difficulty levels.
Figure 8: The distribution of constraint instruction list lengths in datasets of different difficulties.